OpenStudy (anonymous):

What fraction is equivalent to 5%?

OpenStudy (anonymous):

What's another way to write 5%?

OpenStudy (anonymous):

5% is like 1/5

OpenStudy (anonymous):

\[\frac{ 5 }{ 100 }=\frac{ 1 }{ ? }\]

OpenStudy (anonymous):

5/100 or 1/20

OpenStudy (anonymous):

Actually, you don't put it over 100. You put it over 10 but that's okay too :)

OpenStudy (anonymous):

50/100

geerky42 (geerky42):

lol so many wrong answers.

OpenStudy (anonymous):

Hmm, no. Not 50/100 because that's half and 5% is not half.

OpenStudy (anonymous):

10/50

OpenStudy (anonymous):

5% can also be written as .5 because you move the decimal point.

OpenStudy (anonymous):

1/5

geerky42 (geerky42):

you mean .05? @SulcataKid

OpenStudy (anonymous):

Yes, you're right @geerky42 sorry. It's .05 because you move the decimal point twice.

OpenStudy (anonymous):

A. 1/2 B.1/20 C.50/100 D. 10/50

OpenStudy (anonymous):

B

OpenStudy (anonymous):

it is b

geerky42 (geerky42):

AND \(\dfrac{5}{100} \neq \dfrac{1}{25}\)

OpenStudy (anonymous):

I give up. No one wants an explanation. It's B, 1/20

geerky42 (geerky42):

Yep, B...